OpenStudy (anonymous):

Solve : x2 + x + 1/√2 = 0

OpenStudy (dumbcow):

quadratic formula \[x = \frac{-1 \pm \sqrt{1-4/\sqrt{2}}}{2}\]

OpenStudy (anonymous):

@dumbcow can u solve it?

OpenStudy (dumbcow):

i already did

OpenStudy (anonymous):

simplify is joyal. .

OpenStudy (anonymous):

here, , in quadratic equations, there are always 2 roots or value of x, solve it this way post your solution so that you will be guided. .